OpenStudy (coconutjj):

A stone is dropped into some water and a ripple circle of radius r is formed and slowly expands. At the moment the radius is exactly 2m, what rate is the radius of the circle increasing?

OpenStudy (coconutjj):

The perimeter increases at 3m/s, sorry

OpenStudy (lilshane):

what you think

OpenStudy (perl):

A stone is dropped into some water and a ripple circle of radius r is formed and slowly expands. At the moment the radius is exactly 2m, what rate is the radius of the circle increasing? ` The perimeter increases at 3m/s` <--- is this correct?

OpenStudy (coconutjj):

A' = 2pi r

jimthompson5910 (jim_thompson5910):

"The perimeter increases at 3m/s" why not use the circumference formula C = 2*pi*r instead of the area formula

OpenStudy (perl):

yes that would be easier

OpenStudy (perl):

$$ \Large {C =2\pi r \\ \frac{dC}{dt} = 2\pi \cdot \frac{dr}{dt} \\ \therefore \\ 3 = 2\pi \cdot \frac{dr}{dt} \\ \frac{dr}{dt}= \frac{3}{2\pi} } $$ it turns out that it does not matter what is the actual radius

OpenStudy (coconutjj):

sorry internet crashed

OpenStudy (jordan123321):

still need help